OpenStudy (anonymous):

simplify the following expression 5(1-7)+2^3 dvided by (2)(-11)

OpenStudy (anonymous):

1

OpenStudy (anonymous):

woohoo i knew it :)

OpenStudy (anonymous):

5(-6) + 8 divided by -22 = -30+8 all over -22 = 1